Back to the NIKA2 Run 11 main page

Daily Reports

June 13, (Tuesday)

End of the campaign. Here below the preliminary data reduction on two quasars: 1641+399, 2200+420. Stokes I, Q, and U maps are shown. Note that the pattern seen in the Stokes-Q maps of these two quasars is probably due to the instrumental polarization signal. This interpretation will be confirmed by an offline data analysis of unpolarized sources.

1641_I.png 1641_Q.png 1641_U.png

2200_I.png 2200_Q.png 2200_U.png

June 12, (Monday)

14h00 : Observations of Mars start with calibrator on, 23Hz sampling freq, and HWP not in the beam.

Afternoon : Several tests have been performed in order to isolate an additional optical noise we observed when the calibrator is ON (a peak at about 1.3 Hz and additional high frequency noise). See the figures below:

Phase Noise Source OFF - Array 3 (1mm) CalON_00_1mm.png

Phase Noise Source ON - Array 3 (1mm) CalON_11_1mm.png

11h00 : The calibrator and its electronics have been installed and it work nominally.

08h00 : Alain arrived.

05h00 : Poor weather, but Uranus visible, so we start again.

June 11, (Sunday)

Consideration on the NIKA2 Turbo Pump : Looking at the thermometer installed on NIKA2 cryostat (pretty close to the turbo pump) we realised that we had only one night with a strong decrement of the temperature in the receiver cabin (see the next plot). This time corresponds to the night between the 8th and the 9th of June that is to say the only night in which the pump worked continuously until noon of the next day.


18h00 : We stop the observations: very poor weather.

08h00 : Start again. Opacity is still high (around 0.7). Observing Mars.

01h00 : opacity still around 1. Tau225 plot is not updated because linear fit fails (Joaquim dixit) but we can check the exponential fit in

June 10, (Saturday)

18h00 : We stop the observations: very poor weather.

Morning : Observations of Mars at 300mK with a very high tau (about 0.8)

05h30 : Pump on until 07h30 then stop again.

02h05 : sharp raise of temperature from 187mK to 245mk over the next half hour. we did not lose the tuning. We woke up Andrea who turned the pump off. We continue in these conditions.

00h30: The atmosphere and instrument again seem stable. In particular, the histogram of PS amplitudes (in the "kid monitoring" window) is indicative of stable atmosphere conditions, despite relatively high opacities (tau_225 ~0.5).

June 9, (Friday)

Example of noise spectrum obtained with HWP on at nominal configuration (2.98 Hz mechanical frequency), together with the corresponding configuration of the wheel control (to be remembered for future campaigns).


Array 1 Array1_Pol.png

Array 3 Array3_Pol.png

Array 2 Array2_Pol.png

19h30: We start observations with the turbo pump on (1mm arrays at about 200mK). Tau 225 is about 0.5.

17h00: we stop observations because of poor weather.

15h00: Few scans after the pump stopped again. Probably the external temperature is too high to make the temperature control of the receiver cabin working properly. We decided to observe with the pump turned of (1mm arrays at 305mK).

14h00: Observations start after one hour and half with the pump turned on and arrays at about 200mK.

12h30: Turbo Pump stops again.

Morning: Observations of quasars. We did again the characterization of systematics due to the different speeds of the HWP.

05h30: The night run was stable; the instrument and atmosphere (at least via tau_225, and estimate by-eye on the TOD) looked quite stable. However holes in the data have been observed and the acquisition in general seems to have trouble to manage all the instruments at 47 Hz sampling frequency. In particular:

June 8, (Thursday)

In general, very unstable conditions during daytime (mainly anomalous refraction), clearly visible in the quality of the recovered beams (fig. below). Considerable better during the night.


Before midnight, the cryostat looks to be working well, and the array is at ~200mK. We plan to observe a few quasars throughout the night with a "standard" script, with 2.98 Hz HWP rotation. Our plan (for the day) is to observe :

* 18h30: We tried again to turn on the pump. It works now! We start the observations with the turbo pump on set at 30%. The temperature of the 1mm arrays is going to stabilize to about 200mK.

Division of shifts (in local time UT+2h):

10h00: Start of the observations. The plan of the day is to start the observations on a planet (Mars) to characterize the instrumental polarization (i.e intensity to polarization leakage effect). We want also to characterize the noise due to the HWP modulation which is peaked at harmonics of the HWP rotation frequency. Running faster with the HWP means to have access to faster scanning speeds of the telescope but we could have a noise not trivial to subtract from the timelines. In order to find a compromise between this two issues we decided to make tests doing observations with different parameters as follows:

KID Sampling Freq.

HWP Mechanical Rot. Freq.

Tel. Scanning Speed

47.7 Hz

2.98 Hz


47.7 Hz

3.97 Hz


47.7 Hz

1.49 Hz


23.3 Hz

1.49 Hz


* 09h10: Juan turned on the temperature control of the cabin set at 12°C. After 20 minutes, the turbo pump has been switched on at UT set at 30%. It does not work! It switched off several times during the first part of the observations, so we decided to start the observations without it

* 08h30: We measured the noise with the window on the cryostat and a base temperature of 265mK to check if the pulse tube is visible. It does not seem to generate a noise that can bother us during the observations. However, the 1/f noise is large. We hope that after switching on the turbo molecular pump this contribution will be decreased.

Array 1


Array 3


Array 2


June 7, (Wednesday)

* We performed the optical alignment check with Santiago. The alignment is good so we did not change the position of the mirrors. We also checked the alignment of the HWP.

* The electronics of the HWP wheel is working nominally. We tested different speeds without problems.

* Juan Penalver tested the temperature control of the cabin and setting the temperature at 12°C for few hours. Tomorrow before starting the observation we will switch on the turbo pump after setting the temperature of the cabin to 10 or 12°C.

June 6, (Tuesday)

* The temperature of the mixing chamber is about 265 mK without using the turbo molecular pump. The pressure on P101 is slowly increasing showing that the trap is saturating. It does not seems to be a big problem because the slope is about 40mbar/day meaning that probably we can observe until next tuesday without switching the trap.

08h30 Aina, Alessia, Nicolas and Andrea arrived at the telescope. The idea is to prepare carefully the observations starting on Thursday since the last observation in polarisation performed with NIKA2 was more than a year ago. Nicolas Aina and Alessia worked on the polarisation pipeline, Andrea set up the polarisation facilities and check their proper operation.

DailyReportsNika2Run11 (last edited 2017-06-20 09:04:39 by ritaccoa)